# Video Super-Resolution Transformer with Masked Inter&Intra-Frame Attention

### v2

@@ Abstract
Recently, Vision Transformer has achieved great success in recovering missing details in low-resolution sequences, i.e., the video super-resolution (VSR) task. Despite its superiority in VSR accuracy, the heavy computational burden as well as the large memory footprint hinder the deployment of Transformer-based VSR models on constrained devices. In this paper, we address the above issue by proposing a novel feature-level masked processing framework: VSR with Masked Intra and inter frame Attention (MIA-VSR). The core of MIA-VSR is leveraging feature-level temporal continuity between adjacent frames to reduce redundant computations and make more rational use of previously enhanced SR features. Concretely, we propose an intra-frame and inter-frame attention block which takes the respective roles of past features and input features into consideration and only exploits previously enhanced features to provide supplementary information. In addition, an adaptive block-wise mask prediction module is developed to skip unimportant computations according to feature similarity between adjacent frames. We conduct detailed ablation studies to validate our contributions and compare the proposed method with recent state-of-the-art VSR approaches. The experimental results demonstrate that MIA-VSR improves the memory and computation efficiency over state-of-the-art methods, without trading off PSNR accuracy. The code is available at https://github.com/LabShuHangGU/MIA-VSR.
@@ 1 Introduction
In this paper, we provide an affirmative answer to the above question with a novel masked VSR framework, i.e., Masked Inter&Intra frame Attention (MIA) model. Our MIA-VSR method could make use of temporal continuity to recover missing HR details, while, at the same time, leverage temporal redundancy for reducing unnecessary computations. Concretely, our MIA-VSR model advances the existing VSR approaches in the following two aspects. First, we develop a tailored inter-frame and intra-frame attention block (IIA) for making more rational use of previously enhanced features. Instead of directly inheriting the Swin-Transformer block [26] to process concatenated hidden states and image feature, our proposed IIAB block takes into account the respective roles of past features and input features and only utilizes the image feature of the current frame to generate the query token. As a result, the proposed IIAB block not only reduces the computational consumption of the original joint self-attention strategy by a large margin, but also aligns the intermediate feature with spatial coordinates to enable efficient masked processing. Secondly, we propose a feature-level adaptive masked processing mechanism to reduce redundant computations according to the continuity between adjacent frames. Differently from previous efficient processing models, which simply determine skipable regions of the whole network according to pixel intensities, we adopt a feature-level selective skipping strategy and pass by computations of a specific stage adaptively. The proposed feature-level adaptive masking strategy enables our MIA-VSR to save computations and at the same time achieve good VSR results. Our contributions can be summarized as follows.
  • We present a feature-level masked processing framework for efficient VSR, which could take advantage of temporal continuity to reduce redundant computations in the VSR task.
@@
  • We propose an adaptive mask prediction module, which masks out unimportant regions according to feature similarity between adjacent frames for different stages of processing and shows a certain adaptability to test-time adaptation.
  • We compare our MIA-VSR model with state-of-the-art VSR models against which our approach could generate superior results with less computations and memory footprints.
@@ 2.1 Video Super-Resolution
Recurrent framework based VSR. Another category of approaches applies recurrent neural networks to exploit temporal information from more frames. FRVSR [30] first proposed a recurrent framework that utilizes optical flow to align the previous HR estimation and the current LR input for VSR. RLSP [8] propagates high-dimensional hidden states instead of the previous HR estimation to better exploit long-term information. RSDN [12] further extended RLSP [8] by decomposing the LR frames into structure and detail layers and introduced an adaptation module to selectively use the information from hidden state. BasicVSR [3] utilized bi-directional hidden states, and BasicVSR++ [5] further improved BasicVSR with second-order grid propagation and flow-guided deformable alignment. PSRT [31] adopted the bi-directional second-order grid propagation framework of BasicVSR++ and utilized multi-frame self-attention block to jointly process the outputs of the feature propagation of the previous frames and the input features. Generally, by passing the high dimensional hidden states or output feature, the recurrent based VSR could incorporate temporal information from more frames for estimating the missing details and therefore achieve better VSR results. Our proposed MIA-VSR follows the general framework of bi-directional second-order hidden feature propagation while introduces masked processing strategy and intra&inter-frame attention block for the pursuit of better trade-off between VSR accuracy, computational burden and memory footprint.
@@ 2.2 Efficient Video Processing
Various strategies of reducing temporal redundancy have been explored for efficient video processing. A category of approaches adopts the general network optimization strategy and utilizes pruning [36] or distillation [10] methods to train light-weight networks for efficient video processing. In order to take extra benefit from the temporal continuity of video frames, another category of methods only computes expensive backbone features on key-frames and applies concatenation methods [15], optical-flow based alignment methods [40, 19, 41, 15, 28, 11], dynamic convolution methods [28] and self-attention methods [11, 24] to enhance features of other frames with keyframe features. Most recently, [9] proposed a skip-convolution approach which only conducts computation in regions with significant changes between frames to achieve the goal of efficient video processing. However, most of the advanced efficient video processing schemes described above were designed for high-level vision tasks such as object detection and pose estimation. To the best of our knowledge, our study is the first work that leverages temporal continuity across different areas to reduce redundant computation for low-level VSR tasks.
@@ 3 Methodology
Figure 3: Illustration of the inter&intra-frame attention block with adaptive masked processing module. The adaptive mask prediction module (b) in the IIAB block acts in the Attention module’s linear layer which produce the Query, the projection layer and the linear layer in the FFN module during inference to reduce temporal and sptical redundancy calculations (a). $h_{att}$ and $h_{ffn}$ refer to the processed hidden feature in the Attention and FFN module.
@@ 3.1 Overall Architecture
The overall architecture of our proposed MIA-VSR model is shown in Fig.2. We built our MIA-VSR framework upon the bi-directional second-order grid propagation framework of BasicVSR++ [5], which has also been adopted in the recent state-of-the-art method PSRT [31]. The whole model consists of three parts, i.e., the shallow feature extraction part, the recurrent feature refinement part and the feature reconstruction part. We follow previous works BasicVSR++ [5] and PSRT [31] which use a plain convolution operation to extract shallow features and adopt a pixel-shuffle layer [32] to reconstruct HR output with refined features. The patch alignment method used in PSRT [32] is used to align the adjacent frames. We improve the efficiency of existing works by reducing redundant computations in the recurrent feature refinement part.
@@
where $\odot$ is the point-wise multiplication operation. More details of our mask predicting module will be presented in Section 3.3.
@@ 3.2 Inter&Intra-Frame Attention Block
$\left\{\begin{array}[]{l}\bm{K}_{m,n}^{t,intra}=\bm{X}_{m,n}^{t}\bm{W}^{K}_{m,n},\\ \bm{K}_{m,n}^{t,inter}=\left[\bm{X}_{m}^{t-1};\bm{X}_{m}^{t-2}\right]\bm{W}^{K}_{m,n},\par\end{array}\right.$
$\left\{\begin{array}[]{l}\bm{V}_{m,n}^{t,intra}=\bm{X}_{m,n}^{t}\bm{W}^{V}_{m,n},\\ \bm{V}_{m,n}^{t,inter}=\left[\bm{X}_{m}^{t-1};\bm{X}_{m}^{t-2}\right]\bm{W}^{V}_{m,n},\par\end{array}\right.$
@@
where $\bm{K}_{m,n}^{t}=[\bm{K}_{m,n}^{t,inter};\bm{K}_{m,n}^{t,intra}]$ and $\bm{V}_{m,n}^{t}=[\bm{V}_{m,n}^{t,inter};\bm{V}_{m,n}^{t,intra}]$ are the concatenated intra&inter-frame tokens; $d$ is the channel dimension of the Token and $B$ is the learnable relative positional encoding. It should be noted that in Eq.4, all the intermediate inter-frame tokes $\{\bm{V}_{m,n}^{t,inter};\bm{K}_{m,n}^{t,inter}\}_{n=1,\dots,N}$ are generated from the same enhanced features $\left[\bm{x}_{m}^{t-1};\bm{x}_{m}^{t-2}\right]$; which means that we only leverage mature enhanced features from previous frames to provide supplementary information and do not need to utilize the time consuming self-attention mechanism to jointly update the current feature and previous features. An illustration of our proposed inter-frame and intra-frame attention (IIAB) block is presented in Fig.3. In addition to the attention block, our transformer layer also utilizes LayerNorm and FFN layers, which have been commonly utilized in other Transformer-based architectures [21, 31].
@@ 3.3 Adaptive Masked Processing in the Recurrent VSR Transformer
In this subsection, we present details of how we generate block-wise masks $\bm{M}_{m,n}^{t}$ in Eq.3 to further reduce unimportant computations. Habibian et al. [9] first proposed a skip convolution mechanism to reduce redundant computations based on differences between pixel values in adjacent frames. However, for the VSR task, skipping the whole computation process of a certain region according to the intensity differences between adjacent input frames will inevitably affect the accuracy of the model. In this paper, we leverage the temporal continuity to reduce computation in a subtler way: by exploiting feature differences between adjacent frames to select less important computations for each block.
@@ 4.3 Comparison with State-of-the-art Methods
In this subsection, we compare the proposed MIA-VSR model with state-of-the-art VSR methods. We compare the proposed MIA-VSR with representative sliding-window based methods TOFlow [39], EDVR [35], MuCAN [18], VSR-T [2], VRT [20], RVRT [21] and representative recurrent-based methods BasicVSR [3], BasicVSR++ [5], TTVSR [24], PSRT-recurrent [31]; among which VRT, RVRT, TTVSR and PSRT-recurrent are Transformer-based approaches and the other approaches are CNN-based models.
@@
The VSR results by different methods can be found in Table 2. As the number of FLOPs(T) for our MIA-VSR model is related to the content of video sequences, we therefore report the average per frame FLOPs on different datasets. In terms of the results trained with short sequences, the proposed MIA-VSR outperforms the compared methods by a large margin. Our model improves the PSNR of the state-of-the-art PSRT-recurrent model by 0.13 dB with a reduction in the number of FLOPs of almost 40%. As for the models trained with longer training sequences, our MIA-VSR still achieves a better trade-off between VSR accuracy and efficiency over recent state-of-the-art approaches. With more than 40% less computations on the RVRT and PSRT-recurrent approaches, our model achieved the best VSR results on the REDS and Vid4 datasets and the second-best results on the Vimeo-90K dataset. Some visual results from different VSR results can be found in Fig.5, our MIA-VSR method is able to recover more natural and sharp textures from the input LR sequences.
In our supplementary file, we also provide other instantializations of our VSR model to show the scalability of our approach and compare our method with efficient VSR models. More visual examples by different models are also provided.
@@ Appendix B Light-weight MIA-VSR models
Figure 6: PSNR(dB) and FLOPs(G) comparison on the REDS4 [27] dataset. In comparison with the existing video super-resolution methods, our proposed MIA-VSR model, MIA-VSR-small and MIA-VSR-tiny could obtain better trade-offs between VSR results and computational cost. Our fine-tuned model MIA-VSRnormal-$\dagger$ outperforms the current state-of-the-art model by more than 0.1 dB with nearly 40% less number of FLOPs. Our light-weight model MIA-VSR-tiny outperforms the recent light-weight Transformer-based VSR model TTVSR[24] by 0.28 dB, with 45% less number of FLOPs. More details can be found in Section B.

### v4

@@ Abstract
Recently, Vision Transformer has achieved great success in recovering missing details in low-resolution sequences, i.e., the video super-resolution (VSR) task. Despite its superiority in VSR accuracy, the heavy computational burden as well as the large memory footprint hinder the deployment of Transformer-based VSR models on constrained devices. In this paper, we address the above issue by proposing a novel feature-level masked processing framework: VSR with Masked Intra and inter-frame Attention (MIA-VSR). The core of MIA-VSR is leveraging feature-level temporal continuity between adjacent frames to reduce redundant computations and make more rational use of previously enhanced SR features. Concretely, we propose an intra-frame and inter-frame attention block which takes the respective roles of past features and input features into consideration and only exploits previously enhanced features to provide supplementary information. In addition, an adaptive block-wise mask prediction module is developed to skip unimportant computations according to feature similarity between adjacent frames. We conduct detailed ablation studies to validate our contributions and compare the proposed method with recent state-of-the-art VSR approaches. The experimental results demonstrate that MIA-VSR improves the memory and computation efficiency over state-of-the-art methods, without trading off PSNR accuracy. The code is available at https://github.com/LabShuHangGU/MIA-VSR.
@@ 1 Introduction
In this paper, we provide an affirmative answer to the above question with a novel masked VSR framework, i.e., Masked Inter&Intra frame Attention (MIA) model. Our MIA-VSR method could make use of temporal continuity to recover missing HR details, while, at the same time, leveraging temporal redundancy for reducing unnecessary computations. Concretely, our MIA-VSR model advances the existing VSR approaches in the following two aspects. First, we develop a tailored inter-frame and intra-frame attention block (IIAB) for making more rational use of previously enhanced features. Instead of directly inheriting the Swin-Transformer block [27] to process concatenated hidden states and image feature, our proposed IIAB block takes into account the respective roles of past features and input features and only utilizes the image feature of the current frame to generate the query token. As a result, the proposed IIAB block not only reduces the computational consumption of the original joint self-attention strategy by a large margin, but also aligns the intermediate feature with spatial coordinates to enable efficient masked processing. Secondly, we propose a feature-level adaptive masked processing mechanism to reduce redundant computations according to the continuity between adjacent frames. Differently from previous efficient processing models, which simply determine skipable regions of the whole network according to pixel intensities, we adopt a feature-level selective skipping strategy and pass by computations of a specific stage adaptively. The proposed feature-level adaptive masking strategy enables our MIA-VSR to save computations and at the same time achieve good VSR results. Our contributions can be summarized as follows.
  • We present a novel feature-level masked processing framework for efficient VSR, which is able to take advantage of the temporal continuity to reduce redundant computations in the VSR task.
@@
  • We propose an adaptive mask prediction module, which masks out unimportant regions according to the feature similarity between adjacent frames for different stages of processing in the VSR model.
  • We compare our proposed MIA-VSR model with state-of-the-art VSR models against which our approach could generate superior results with less computations and memory footprints.
@@ 2.1 Video Super-Resolution
Recurrent framework based VSR. Another category of approaches applies recurrent neural networks to exploit temporal information from more frames. FRVSR [31] first proposed a recurrent framework that utilizes optical flow to align the previous HR estimation and the current LR input for VSR. RLSP [8] propagates high-dimensional hidden states instead of the previous HR estimation to better exploit long-term information. RSDN [13] further extended RLSP [8] by decomposing the LR frames into structure and detail layers and introduced an adaptation module to selectively use the information from hidden state. BasicVSR [3] utilized bi-directional hidden states, and BasicVSR++ [5] further improved BasicVSR with second-order grid propagation and flow-guided deformable alignment. PSRT [32] adopted the bi-directional second-order grid propagation framework of BasicVSR++ and utilized multi-frame self-attention block to jointly process the outputs of the feature propagation of the previous frames and the input features. Generally, by passing the high dimensional hidden states or output feature, the recurrent based VSR could incorporate temporal information from more frames for estimating the missing details and therefore achieve better VSR results. Our proposed MIA-VSR follows the general framework of bi-directional second-order hidden feature propagation while introducing masked processing strategy and intra&inter-frame attention block for the pursuit of better trade-off between VSR accuracy, computational burden and memory footprint.
@@ 2.2 Efficient Video Processing
Various strategies of reducing temporal redundancy have been explored for efficient video processing. A category of approaches adopts the general network optimization strategy and utilizes pruning [37] or distillation [11] methods to train light-weight networks for efficient video processing. In order to take extra benefit from the temporal continuity of video frames, another category of methods only computes expensive backbone features on key-frames and applies concatenation methods [16], optical-flow based alignment methods [41, 20, 42, 16, 29, 12], dynamic convolution methods [29] and self-attention methods [12, 25] to enhance features of other frames with keyframe features. Most recently, [10] proposed a skip-convolution approach which only conducts computation in regions with significant changes between frames to achieve the goal of efficient video processing. However, most of the advanced efficient video processing schemes described above were designed for high-level vision tasks such as object detection and pose estimation. To the best of our knowledge, our study is the first work that leverages the temporal continuity across different areas for each block to reduce redundant computation for low-level VSR tasks.
@@ 3 Methodology
Figure 3: Illustration of the inter&intra-frame attention block with adaptive masked processing module. The adaptive mask prediction module (b) in the IIAB block acts in the Attention module’s linear layer which produce the Query, the projection layer and the linear layer in the FFN module during inference to reduce temporal and sptical redundancy calculations (a). ${\bm{X}_{m,n}^{t}}^{\prime}$ and ${\bm{X}_{m,n}^{t}}^{\prime\prime}$ refer to the processed hidden feature in the Attention and FFN module.
@@ 3.1 Overall Architecture
The overall architecture of our proposed MIA-VSR model is shown in Fig.2. We built our MIA-VSR framework upon the bi-directional second-order grid propagation framework of BasicVSR++ [5], which has also been adopted in the recent state-of-the-art method PSRT [32]. The whole model consists of three parts, i.e., the shallow feature extraction part, the recurrent feature refinement part and the feature reconstruction part. We follow previous works BasicVSR++ [5] and PSRT [32] which use a plain convolution operation to extract shallow features and adopt a pixel-shuffle layer [33, 9] to reconstruct HR output with refined features. The patch alignment method used in PSRT [33] is used to align the adjacent frames. We improve the efficiency of existing works by reducing redundant computations in the recurrent feature refinement part.
@@
where $\bm{\hat{X}}^{t}_{m,n}$ is the processed hidden feature in the IIAB block (shown as ${\bm{X}_{m,n}^{t}}^{\prime}$ and ${\bm{X}_{m,n}^{t}}^{\prime\prime}$ in Fig.3) and $\odot$ is the point-wise multiplication operation. More details of our mask predicting module will be presented in Section 3.3.
@@ 3.2 Inter&Intra-Frame Attention Block
$\left\{\begin{array}[]{l}\bm{K}_{m,n}^{t,intra}=\bm{X}_{m,n}^{t}\bm{W}^{K}_{m,n},\\ \bm{K}_{m,n}^{t,inter}=\left[\bm{X}_{m+1}^{t-1};\bm{X}_{m+1}^{t-2}\right]\bm{W}^{K}_{m,n},\end{array}\right.$
$\left\{\begin{array}[]{l}\bm{V}_{m,n}^{t,intra}=\bm{X}_{m,n}^{t}\bm{W}^{V}_{m,n},\\ \bm{V}_{m,n}^{t,inter}=\left[\bm{X}_{m+1}^{t-1};\bm{X}_{m+1}^{t-2}\right]\bm{W}^{V}_{m,n},\end{array}\right.$
@@
where $\bm{K}_{m,n}^{t}=[\bm{K}_{m,n}^{t,inter};\bm{K}_{m,n}^{t,intra}]$ and $\bm{V}_{m,n}^{t}=[\bm{V}_{m,n}^{t,inter};\bm{V}_{m,n}^{t,intra}]$ are the concatenated intra&inter-frame tokens; $d$ is the channel dimension of the Token and $B$ is the learnable relative positional encoding. It should be noted that in Eq.4, all the intermediate inter-frame tokes $\{\bm{V}_{m,n}^{t,inter};\bm{K}_{m,n}^{t,inter}\}_{n=1,\dots,N}$ are generated from the same enhanced features $\left[\bm{X}_{m+1}^{t-1};\bm{X}_{m+1}^{t-2}\right]$; which means that we only leverage mature enhanced features from previous frames to provide supplementary information and do not need to utilize the time consuming self-attention mechanism to jointly update the current feature and previous features. An illustration of our proposed inter-frame and intra-frame attention (IIAB) block is presented in Fig.3. In addition to the attention block, our proposed transformer layer also utilizes LayerNorm and FFN layers, which have been commonly utilized in other Transformer-based VSR architectures [22, 32].
@@ 3.3 Adaptive Masked Processing in the Recurrent VSR Transformer
In this subsection, we present details of how we generate block-wise masks $\bm{M}_{m,n}^{t}$ in Eq.3 to further reduce unimportant computations. Habibian et al. [10] first proposed a skip convolution mechanism to reduce redundant computations based on differences between pixel values in adjacent frames. However, for the VSR task, skipping the whole computation process of a certain region according to the intensity differences between adjacent input frames will inevitably affect the accuracy of the model. In this paper, we leverage the temporal continuity to reduce the computation of the VSR model in a subtler way: by exploiting feature differences between adjacent frames to select less important computations for each block.
@@ 4.3 Comparison with State-of-the-art Methods
In this subsection, we compare the proposed MIA-VSR model with state-of-the-art VSR methods. We compare the MIA-VSR with representative sliding-window based methods TOFlow [40], EDVR [36], MuCAN [19], VSR-T [2], VRT [21], RVRT [22] and representative recurrent-based methods BasicVSR [3], BasicVSR++ [5], TTVSR [25], PSRT-recurrent [32]; among which VRT, RVRT, TTVSR and PSRT-recurrent are Transformer-based approaches and the other approaches are CNN-based models.
@@
The VSR results by different methods can be found in Table 2. As the number of FLOPs(T) for our MIA-VSR model is related to the content of video sequences, we therefore report the average per frame FLOPs on different datasets. In terms of the results trained with short sequences, the proposed MIA-VSR outperforms the compared methods by a large margin. Our model improves the PSNR of the state-of-the-art PSRT-recurrent model by 0.13 dB with a reduction in the number of FLOPs of almost 40%. As for the models trained with longer training sequences, our MIA-VSR still achieves a better trade-off between VSR accuracy and efficiency over recent state-of-the-art approaches. With more than 40% less computations on the RVRT and PSRT-recurrent approaches, our model achieved the best VSR results on the REDS and Vid4 datasets and the second-best results on the Vimeo-90K dataset. Some visual results from different VSR results can be found in Fig.5, our proposed MIA-VSR method is able to recover more natural and sharp textures from the input LR video sequences.
In our supplementary file, we also provide other instantializations of our proposed MIA-VSR model to show the scalability of our approach and compare our method with efficient VSR models. More visual examples are also provided for different datasets.
@@ Appendix B Light-weight MIA-VSR models
[…]